### v1

# Permutation-Invariant and Orientation-Aware Dataset Distillation for 3D Point Clouds

Abstract
We should collect large amount of data to train deep neural networks for various applications. Recently, the dataset distillation for images and texts has been attracting a lot of attention, that reduces the original dataset to a synthetic dataset while preserving essential task-relevant information. However, 3D point clouds distillation is almost unexplored due to the challenges of unordered structures of points. In this paper, we propose a novel distribution matching-based dataset distillation method for 3D point clouds that jointly optimizes the geometric structures of synthetic dataset as well as the orientations of synthetic models. To ensure the consistent feature alignment between different 3D point cloud models, we devise a permutation invariant distribution matching loss with the sorted feature vectors. We also employ learnable rotation angles to transform each syntheic model according to the optimal orientation best representing the original feature distribution. Extensive experimental results on widely used four benchmark datasets, including ModelNet10, ModelNet40, ShapeNet, and ScanObjectNN, demonstrate that the proposed method consistently outperforms the existing methods.
1 Introduction
With the increasing demand for large-scale datasets, dataset distillation [21] has gained significant attention across various applications of deep learning. The goal of dataset distillation is to reduce the size of an original dataset into a synthetic dataset that preserves the essential task-relevant information of the original dataset. Therefore, the models trained on the reduced synthetic dataset are encouraged to achieve comparable performance to those trained on the original dataset. Existing dataset distillation methods [3, 6, 8, 11, 29, 28, 30, 17, 26, 5, 9] can be broadly classified into gradient matching, trajectory matching, and distribution matching approaches. The gradient matching and trajectory matching methods aim to ensure that the synthetic dataset produces similar optimization dynamics to the original dataset. To this end, the gradient matching method [29] minimizes the difference between the gradients computed on the synthetic and original datasets. However, these methods potentially overlook long-term dependencies in the training process. Rather than comparing individual gradients, the trajectory matching methods [3, 6, 8, 11] encourage the models trained on a synthetic dataset follow similar optimization trajectories to those trained on the original dataset. However, these methods train the networks while optimizing the synthetic dataset significantly increasing the computational complexity. To alleviate the computational burden in dataset distillation, the distribution matching techniques [28, 30, 17, 26, 5] have been introduced. They randomly initialize the networks without training to extract the features, and compare the feature distributions between the original and synthetic datasets.
Figure 1: The concept of the proposed dataset distillation method for 3D point clouds. (a) Permutation-invariant feature mathcing. (b) Orientation optimization.
While the dataset distillation has been extensively studied for structured data, such as images [29, 3, 28] and texts [10, 12, 13], that for 3D point clouds is almost unexplored. Recently, PCC [27] applied an existing gradient matching method of DC [29] to tackle the data distillation problem of 3D point clouds, but it did not consider the inherent characteristics of 3D point clouds and still suffers from the high computational cost. Unlike structured data, 3D point clouds consist of unordered and irregularly distributed points in 3D space. Specifically, inconsistent indexing of points makes it challenging to develop efficient dataset distillation techniques. Furthermore, 3D models often exhibit different orientation from one another, that also degrades the distillation performance.
In this paper, we propose a joint optimization framework of Permutation Invariant Distribution Matching (PIDM) and orientation optimization for dataset disllation of 3D point clouds. A major challenge in this setting is achieving effective feature alignment despite the unordered nature of point clouds and their arbitrary orientations. As illustrated in Figure 1(a), the PIDM loss addresses the issue of unordered structure by sorting the elements within each feature vector according to their sizes before computing the distance between two feature distributions. To further enhance the feature alignment, as shown in Figure 1(b), We also employ learnable parameters of the rotation angles to estimate optimal orientation of 3D models, while optimizing the synthetic dataset. Experimental results demonstrate that the proposed method optimizes both the synthetic dataset and the orientation of 3D models simultaneously outperforming the existing methodologies.
The key contributions of this paper are summarized as follows:
  • To the best of our knowledge, we are the first to proposed a distribution matching-based dataset distillation method for 3D point clouds, that jointly optimizes the synthetic dataset and the orientation of 3D models simultaneously.
  • We devised PIDM loss to ensure permutation invariant feature alignment for 3D point clouds with inconsistent orders of points.
  • We validated proposed method through comprehensive experiments on four benchmark datasets of 3D point clouds classification, and showed the superiority of the proposed method over the existing dataset distillation techniques.
2 Related Works
Figure 2: The overall framework of the proposed 3D point clouds dataset distillation method.
2.1 3D Point Data Analysis
3D point clouds are generally unordered exhibiting irregular characteristics, that makes it difficult to apply the convolution operations in deep neural networks commonly used for images. Early approaches [25, 32] convert the point clouds into structured representation of multi-view images or voxel grids to enable the use of conventional deep learning architectures, such as convolutional neural networks (CNNs). VoxelNet [32] voxelizes the point clouds, but it causes increased memory consumption and quantization errors. To directly process point clouds without conversion to intermediate structured representations, PointNet [14] was introduced as a pioneering work that learns the features directly from unordered raw point clouds. However, it suffers from the limitation in capturing local geometric relationships, and hierarchical CNNs have been developed tailored to 3D point clouds. PointNet++[15] captures local geometric structures through iterative sampling and grouping strategies. On the other hand, PointConv [23] generalizes the convolution operations to irregular point clouds by adaptively weighting the neighboring points based on their spatial distribution. Meanwhile, DGCNN [22] constructs dynamic graphs from point clouds to adaptively capture the semantic information. Attempts have been also made to apply the transformer [20] to 3D point clouds processing, where the Point Transformer [31] utilizes the attention mechanism to capture the long-range dependencies.
2.2 Coreset Selection
Coreset selection is a technique designed to select representative samples from a given dataset, while maintaining the model’s performance even with the sampled data. The random selection method [16] randomly chooses a subset of data samples from the whole dataset. It is simple but suffers from the robustness due to the lack of informative criteria for sampling. The K-center method [18] selects data points iteratively that maximize the minimum distance to the set of already selected ones, taking into account the data distribution. The Herding method [2, 1] iteratively takes data points that minimize the discrepancy between the mean embeddings of the selected subset and the entire dataset in the feature space.
2.3 Dataset Distillation
Dataset distillation [21] methods can be largely categorized into the gradient matching [29], trajectory matching [3, 6, 8, 11], and distribution matching [28, 30, 17, 26, 5] approaches. The gradient matching, first introduced by DC [29], minimizes the difference between the gradients computed from the original and synthetic datasets, respectively, guiding the synthetic dataset to follow the training direction of the original dataset. The trajectory matching methods [3, 6, 8, 11], initially proposed by MTT [3], encourage the models trained on the synthetic dataset to follow the optimization trajectories similar to those trained on the original dataset rather than comparing individual gradients. ATT [11] automatically adjusts the length of the training trajectories between the synthetic and original datasets, enabling more effective and precise matching. The distribution matching, introduced by DM [28], focuses on minimizing the distance between the feature distributions of the original and synthetic datasets. DataDAM [17] enhances the distribution matching by aligning the feature maps using attention mechanism, achieving unbiased feature representation with low computational overhead. Furthermore, M3D [26] employs the Gaussian kernel function in the distribution matching loss, enabling the alignment across higher-order statistical characteristics of the feature distributions. Recently, a gradient matching-based point clouds distillation method [27] has been introduced, however it simply applied the existing image-based method without considering the unique characteristics of 3D point clouds.
3 Methodology
We propose a novel dataset distillation method for 3D point clouds that first addressing the key challenges of unordered structure and rotational variation of 3D point clouds. We perform PIDM that effectively aligns the features of unordered points across different 3D models. To handle the rotational variation, we also estimate optimal orientation while generating the synthetic dataset. Figure 2 shows the overall framework of the proposed method.
3.1 Problem Formulation
The objective of dataset distillation is to compress the task-relavant information in the original dataset ${\boldsymbol{\mathcal{T}}}=\left\{{{\bf t}}_{i}\right\}$ and generate a much smaller synthetic dataset ${\boldsymbol{\mathcal{S}}}=\left\{{{\bf s}}_{i}\right\}$, $|{\boldsymbol{\mathcal{S}}}|\ll|{\boldsymbol{\mathcal{T}}}|$, where the models trained on ${\boldsymbol{\mathcal{S}}}$ achieve close performance to that trained on ${\boldsymbol{\mathcal{T}}}$. Given a 3D point cloud sample ${{\bf p}}$ following a real data distribution with the corresponding class label $l$, the optimal synthetic dataset ${\boldsymbol{\mathcal{S}}}^{\star}$ can be obtained via
${\boldsymbol{\mathcal{S}}}^{\star}=\mathop{\arg\min}\limits_{{\boldsymbol{\mathcal{S}}}}\mathbb{E}_{{{\bf p}}}\left[||{\mathcal{L}}(\phi_{{{\boldsymbol{\mathcal{T}}}}}({{\bf p}}),l)-{\mathcal{L}}(\phi_{{{\boldsymbol{\mathcal{S}}}}}({{\bf p}}),l)||^{2}\right],$ (1)
where ${\mathcal{L}}$ denotes a task-specific loss function, such as the cross-entropy loss, and $\phi_{{\boldsymbol{\mathcal{T}}}}$ and $\phi_{{\boldsymbol{\mathcal{S}}}}$ represent the models trained on ${\boldsymbol{\mathcal{T}}}$ and ${\boldsymbol{\mathcal{S}}}$, respectively.
3.2 Permutation Invariant Distribution Matching
We basically perform the distribution matching by minimizing the discrepancy between the feature distributions derived from the original and synthetic datasets, respectively, such that
${\boldsymbol{\mathcal{S}}}^{\star}=\mathop{\arg\min}_{{\boldsymbol{\mathcal{S}}}}D(\phi({\boldsymbol{\mathcal{T}}}),\phi({\boldsymbol{\mathcal{S}}})),$ (2)
where $D$ is a distance function, and $\phi$ denotes a randomly initialized untrained network. Note that we simply use a randomly initialized network without training and effectively alleviate the computational overhead in dataset distillation. The Maximum Mean Discrepancy (MMD) [7] loss $\mathcal{L}_{\text{MMD}}$ is often used to measure $D$, defined as
$\mathcal{L}_{\text{MMD}}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}})={K}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{T}}})+{K}({\boldsymbol{\mathcal{S}},\boldsymbol{\mathcal{S}}})-2{K}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}}),$ (3)
where ${K}(\cdot,\cdot)$ is the kernel function, where we used the Gaussian kernel in this work, given by
${K}({{\boldsymbol{\mathcal{T}}},\boldsymbol{\mathcal{S}}})=\frac{1}{|\boldsymbol{\mathcal{T}}|\cdot|\boldsymbol{\mathcal{S}}|}\sum_{{{{\bf t}}\in\boldsymbol{\mathcal{T}}}}\sum_{{{\bf s}}\in\boldsymbol{\mathcal{S}}}\text{exp}\left(-\frac{||\phi({{\bf t}})-\phi({{\bf s}})||^{2}}{2\sigma}\right).$ (4)
Note that the spatial relationships among the pixels are well structured and consistent across different images, but the points in 3D space are usually indexed in different orders across multiple 3D models. Therefore, the conventional distribution matching methods cannot be directly applied to 3D point clouds, since they assume aligned features between two compared models. To overcome this issue, we propose PIDM that performs the feature sorting in advance when comparing the feature distributions between the original and synthetic datasets. Specifically, given a 3D point cloud model ${{\bf p}}$, we extract point-wise features by using $\phi$ as
$\phi({{\bf p}})=\left[{\bf f}_{1},{\bf f}_{2},\dots,{\bf f}_{C}\right],\quad{\bf f}_{i}\in\mathbb{R}^{N},$ (5)
where $N$ denotes the number of points in ${{\bf p}}$, and $C$ is the number of feature channels. We first rearrange the elements of ${\bf f}_{i}$ in descending order of their sizes to have the sorted feature vector $\tilde{{\bf f}}_{i}$. Then we have the sorted features as
$\tilde{\phi}({{\bf p}})=\left[\tilde{{\bf f}}_{1},\tilde{{\bf f}}_{2},\dots,\tilde{{\bf f}}_{C}\right],\quad\tilde{{\bf f}}_{i}\in\mathbb{R}^{N}.$ (6)
Note that ${{\bf t}}$ in the original dataset $\boldsymbol{\mathcal{T}}$ and ${{\bf s}}$ in the synthetic dataset $\boldsymbol{\mathcal{S}}$ exhibit different orders of points in general, and therefore the features in ${\phi}({{\bf t}})$ and ${\phi}({{\bf s}})$ are inconsistenly aligned with each other disrupting the desired behavior of comparison in (4). On the contrary, $\tilde{\phi}({{\bf t}})$ and $\tilde{\phi}({{\bf s}})$ yield well aligned features invariant to the permutation variation, and therefore, we compute the refined Gaussian kernel with the sorted features, given by
$\tilde{K}({{\boldsymbol{\mathcal{T}}},\boldsymbol{\mathcal{S}}})=\frac{1}{|\boldsymbol{\mathcal{T}}|\cdot|\boldsymbol{\mathcal{S}}|}\sum_{{{{\bf t}}\in\boldsymbol{\mathcal{T}}}}\sum_{{{\bf s}}\in\boldsymbol{\mathcal{S}}}\text{exp}\left(-\frac{||\tilde{\phi}({{\bf t}})-\tilde{\phi}({{\bf s}})||^{2}}{2\sigma}\right).$ (7)
Accordingly, we devise the PIDM loss $\mathcal{L}_{\text{PIDM}}$ from $\mathcal{L}_{\text{MMD}}$, given by
$\mathcal{L}_{\text{PIDM}}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}})={K}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{T}}})+{K}({\boldsymbol{\mathcal{S}},\boldsymbol{\mathcal{S}}})-2\tilde{K}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}}).$ (8)
By using $\mathcal{L}_{\text{PIDM}}$ instead of $\mathcal{L}_{\text{MMD}}$, we improve the performance of datasets distillation for 3D point clouds regardless of the orders of points.
Datasets ModelNet10 [24] ModelNet40 [24] ShapeNet [4] ScanObjectNN [19]
PPC 1 3 10 1 3 10 1 3 10 1 3 10
Ratio (%) 0.25 0.75 2.5 0.4 1.2 4.0 0.15 0.45 1.5 0.15 0.45 1.5
Whole 91.41 87.84 82.49 63.84
Random 35.52 75.17 85.31 34.57 59.96 74.06 33.96 54.84 63.13 13.85 20.42 34.75
Herding 40.07 78.02 86.86 54.38 68.45 78.84 49.49 59.79 66.85 15.67 27.66 38.67
K-center 40.07 77.59 83.22 54.38 63.03 65.26 49.49 51.39 47.82 15.67 19.81 24.00
DM 31.86 77.56 86.10 32.29 61.97 75.23 27.84 56.06 64.27 14.89 23.25 35.58
DC 43.89 74.97 86.11 52.00 66.56 75.57 49.78 60.08 64.58 18.63 24.30 35.27
Ours 44.70 84.96 87.79 55.80 72.08 80.07 50.20 63.74 68.35 17.29 31.84 43.91
Table 1: Performance comparison between three coreset selection methods [16, 2, 18], DC [29], DM [28], and the proposed method. DC, DM, and proposed method were initialized with random selection [16] for fair comparison. The best and the second best scores are highlighted in bold and underlined, respectively.
3.3 Orientation Optimization
3D models usually exhibit different orientations with each other. Therefore, while generating optimal 3D models in the synthetic dataset in terms of the geometric shapes, we also estimate their optimal orientations best representing the various orientations of 3D models in the original dataset. In practice, we employ three rotation angles of $\theta_{x}$, $\theta_{y}$, and $\theta_{z}$, around the $x$, $y$, and $z$-axes, respectively, as learnable parameters which are used to adaptively adjust the orientation of each 3D model in the synthetic dataset ${\boldsymbol{\mathcal{S}}}$ during the dataset distillation. Therefore, instead of $\mathcal{L}_{\text{PIDM}}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}})$ in (8), we minimize $\mathcal{L}_{\text{PIDM}}({\boldsymbol{\mathcal{T}},\hat{\boldsymbol{\mathcal{S}}}})$, where $\hat{\boldsymbol{\mathcal{S}}}$ is the synthetic dataset where its 3D models are rotated toward the estimated optimal orientation.
Note that we optimize the synthetic 3D models as well as their orientations simultaneously, during the dataset distillation process. Specifically, at each iteration, we randomly initialize the network parameters and construct the synthetic dataset $\boldsymbol{\mathcal{S}}$ by randomly selecting samples from the original dataset. We then have $\hat{\boldsymbol{\mathcal{S}}}$ by rotating the 3D models in $\boldsymbol{\mathcal{S}}$ according to the angles $\theta_{x}$, $\theta_{y}$, and $\theta_{z}$, initially set to 0. Then, for each class, a mini-batch is sampled from the original dataset $\boldsymbol{\mathcal{T}}$, with a batch size of 8 per each class. The corresponding synthetic mini-batch is sampled from $\hat{\boldsymbol{\mathcal{S}}}$, with its batch size set to the number of point clouds per class (PPC). The joint optimization process iteratively updates both the geometric structure of the synthetic dataset and the orientation parameters of its models by minimizing $\mathcal{L}_{\text{PIDM}}({\boldsymbol{\mathcal{T}},\hat{\boldsymbol{\mathcal{S}}}})$. This ensures that the synthetic dataset preserves the geometric characteristics of the original dataset while optimizing the orientation.
4 Experimental Results
4.1 Experimental Setup
Datasets. The proposed method was evaluated on the ModelNet10 [24], ModelNet40 [24], ShapeNet [4], and ScanObjectNN [19] datasets. The ModelNet10, ModelNet40, and ShapeNet are synthetic datasets generated from CAD models, containing 10, 40, and 55 classes, respectively. The ScanObjectNN consists of real-world scanned objects from 15 classes, and hence a more challenging dataset due to the presence of noise, occlusions, and other artifacts commonly found in scanned 3D models.
Evaluation Metrics. Following previous methods [29, 28], we measure the classification accuracy trained on the distilled synthetic dataset. Since the training dynamics of the network can vary with each run, we ensure the fairness by training the network 10 times and report the mean accuracy across these runs. Similar to Images Per Class (IPC) in the previous image-based distillation, we evaluate the performance across different Point clouds Per Class (PPC) values. Experiments were conducted under the PPC settings of 1, 3, and 10 to analyze the impact of varying data compression levels.
Implementation Details. Each point cloud model contains 1,024 points, which is a common standard used in the 3D point clouds classification tasks [14, 15]. The dataset distillation process for the synthetic dataset was optimized for 1,500 iterations, where the performance was evaluated at every 250 iteration by training the PointNet on the synthetic dataset and testing it on the original test dataset. Additionally, SGD was used as the optimizer for $\boldsymbol{\mathcal{S}}$, with a learning rate of 10. For optimizing the rotation angles, SGD was also used, with a learning rate of 0.5 for $\theta_{x}$ and $\theta_{z}$. Since the most datasets are aligned with the y-axis, a higher learning rate of 5 was applied for optimizing $\theta_{y}$. PointNet was optimized using the SGD optimizer with a learning rate of 0.01, and the training was conducted for 500 iterations. Also, note that we only use the distribution of the features before the max pooling to compute the PIDM loss to retain fine-grained structural information. To further emphasize the most dominant features after sorting, the loss computed from the largest magnitude values is also used.
4.2 Performance Comparison
We compare the three most representative coreset selection methods: random selection [16], Herding [2, 1], and K-Center [18]. We additionally compare DC [29] and DM [28], representative distillation methods for 2D images, to provide a broader perspective on dataset distillation. We also compare PCC [27], the only existing 3D point clouds dataset distillation method.
Table 1 compares the quantitative performance of the proposed method with that of the existing methodologies. DM and DC were initialized with random selection [16]. The synthetic datasets were optimized using PointNet [14], and classification performance was evaluated on PointNet to demonstrate their validity. Among the coreset selection methods, Herding [2, 1] achieves the best overall performance. When PPC is set to 1, Herding and K-Center [18] yield identical results since K-Center selects the first data point using the same algorithm to Herding. As PPC increases, Herding consistently outperforms other coreset selection methods, reflecting its effectiveness. However, since coreset selection methods rely only on sample selection without further optimization, they struggle to capture complex feature distributions.
DM [28] matches the distribution of the global features extracted by PointNet. However, PointNet applies the global max pooling before classification, retaining only the most dominant signal per channel while discarding substantial information. This loss of information weakens the effectiveness of the distribution matching, preventing DM from accurately capturing the feature distribution of the original dataset. DC [29], on the other hand, tends to fit the specific network used for distillation due to its structure designed for learning the network. However, PointNet [14], which is used for distillation, is significantly larger than typical networks in image-based distillation tasks, making it difficult for DC to precisely align the gradients. As a result, it consistently underperforms compared to the proposed method. Furthermore, as the PPC increases, the optimization process involves a larger number of point clouds, making the gradient alignment increasingly unstable and leading to reduced performance.
The proposed method outperforms both the coreset selection and existing distillation techniques. When the PPC is small, accurately capturing the original distribution is challenging, leading less noticeable improvement. As the PPC increases, the method more effectively matches the feature distributions significantly improving the performance.
Datasets ModelNet10 [24] ModelNet40 [24] ShapeNet [4] ScanObjectNN [19]
Method DM DC Ours DM DC Ours DM DC Ours DM DC Ours
PointNet++ 34.97 51.49 82.44 14.14 46.88 68.45 20.26 34.30 53.20 15.45 11.57 19.70
DGCNN 56.60 62.58 79.01 51.65 52.30 66.86 13.77 37.98 52.81 17.72 12.49 19.02
PointConv 33.41 39.96 56.57 20.00 37.90 51.32 17.35 33.99 47.41 15.16 14.20 16.82
PT 48.52 61.81 77.62 26.70 47.54 61.95 34.15 42.62 55.48 17.06 14.89 21.87
Table 2: Comparison of cross-architecture generalization performance of DC [29], DM [28], and the proposed method, evaluated on PointNet++ [15], DGCNN [22], PointConv [23], and PT [31], respectively.
Init Method MN10 MN40 SN SONN
Noise DC 18.77 6.05 5.24 12.73
DM 26.53 12.04 12.09 13.55
Ours 70.69 65.25 58.44 29.03
Random DC 68.32 64.71 58.15 26.07
DM 65.17 56.50 49.39 24.57
Ours 72.48 69.31 60.76 31.01
K-center DC 69.57 61.85 42.88 21.65
DM 63.04 49.93 44.82 21.59
Ours 72.72 68.20 60.70 28.25
Herding DC 71.50 66.79 59.23 26.98
DM 64.76 59.16 52.61 27.58
Ours 72.93 69.67 62.25 31.08
Table 3: Performance comparison across different initialization strategies including uniform noise, random selection [16], Herding [2], and K-center [18]. We averaged the accuracies of thre cases of PPC, 1, 3, and 10. MN10: ModelNet10. MN40: ModelNet40. SN: ShapeNet. SONN: ScanObjectNN.
4.2.1 Cross-Architecture Generalization
We also present the cross-architecture generalization performance in Table 2, where the synthetic dataset is optimized using PointNet [14] and evaluated on different backbone networks including PointNet++[15], DGCNN[22], PointConv [23], and PT [31]. Unlike PointNet, the architectures used for generalization performance comparison follow a hierarchical structure similar to CNN-based networks in image processing, exhibiting significantly different characteristics from PointNet. This difference makes it essential for a distilled dataset to retain rich and transferable features that generalize well across architectures.
Since the use of symmetric functions like global max pooling causes significant information loss, DM fails to preserve the original feature distribution and instead learns only the features relevant to PointNet. This limitation prevents the synthetic dataset from capturing diverse structural details, making it highly specific to PointNet. As a result, when evaluated on hierarchical models, DM performs poorly, as it lacks the necessary information for effective generalization. DC, while performing better than DM, also suffers from key limitations. Its gradient matching approach forces the synthetic dataset to align with whole gradients of PointNet parameters, which makes it more effective within the same architecture but limits its adaptability to other models. Also, the bi-level optimization process further aggravates this issue by overfitting the synthetic dataset to the specific training conditions of PointNet. Consequently, when applied to other architectures, DC struggles to generalize effectively.
The proposed method addresses these limitations by direct feature distribution matching with randomly initialized network instead of bi-level optimization. By comparing feature maps with richer information, the synthetic dataset captures features relevant to PointNet while following the original data distribution more closely. This broader feature matching strategy improves performance across different architectures. Unlike DC, which suffers from overfitting, and DM, which loses critical information, the proposed approach adapts more effectively, showing strong generalization across various backbone networks.
Method MN10 MN40 SN SONN
DC 1.52 5.43 7.57 2.19
DM 0.04 0.11 0.15 0.05
Ours 0.08 0.26 0.35 0.11
Table 4: Comparison of training time (in hours) of DC, DM, and the proposed method. Experiments were conducted at PPC 3. MN10: ModelNet10. MN40: ModelNet40. SN: ShapeNet. SONN: ScanObjectNN.
Figure 3: Comparison of synthetic datasets distilled from the ModelNet40 [24] dataset. Point clouds were colorized according to the y-coordinates. From top to bottom, the orginal dataset (first), initialized point cloud models (second), and the distilled synthetic datasets by using DC [29] (third), DM [28] (fourth), and the proposed method (fifth).
MN10 MN40 SN SONN
Random 75.17 59.96 54.84 20.42
w/o sorting 75.01 59.96 53.91 20.20
w/ sorting 84.96 72.08 63.74 31.57
Table 5: Ablation study of the proposed sorting method. Experiments were performed at PPC 3. “Random” denotes the results of the random selection [16]. MN10: ModelNet10. MN40: ModelNet40. SN: ShapeNet. SONN: ScanObjectNN.
4.2.2 Initialization Strategies
Table 3 presents the classification performance on PointNet for DC [29], PCC [27] and DM [28] under different initialization strategies, including uniform noise, random selection [16], Herding [2, 1], and K-Center [18], compared to the proposed method. Note that the results of PCC are obtained from our implementation, where PCC refers to performing DC on a dataset initialized using Herding.
When initialized with noise, DC fails to effectively optimize the synthetic dataset, as the lack of geometric information disrupts the gradient computation. As a result, the original dataset fails to guide the training process along a meaningful path, leading to poor convergence. In contrast, structured initializations like random selection, Herding, or K-Center improve performance, with Herding performing best by selecting the most representative samples. Unlike DC, DM does not include network training process, making it less sensitive to noise initialization and allowing for better performance in such cases. However, when initialized with random selection, K-Center, or Herding, DM consistently underperforms DC due to its reliance on feature representations that discard significant information.
Unlike DC or DM, whose dataset distillation algorithms are unstable, the proposed method achieves comparable performance even when initialized from uniform noise. Furthermore, the proposed method significantly outperforms other approaches under the heuristic initialization strategy, such as random selection. This demonstrates that the proposed method is robust to initialization and effectively resolves the instability issues inherent in previous dataset distillation approaches.
${\boldsymbol{\mathcal{S}}}$ ${\boldsymbol{\mathcal{\theta}}}$ Aligned Mixed Rotated
_ _ 65.01 59.96 20.42
_ 66.22 62.17 24.89
_ 73.96 71.51 29.72
74.35 72.08 31.84
Table 6: Ablation study of optimizing the synthetic dataset and orientation. Experiments were performed at PPC 3. The baseline (first row) shows the performance without optimization which is identical to that of the random selection.
4.2.3 Training Time Comparison
Table 4 compares the training time between the proposed and previous methods. DM [28] achieves the fastest training time as it uses only a small subset of the overall features. However, this limited feature usage causes many useful features to be ignored, leading to performance degradation. In contrast, DC [29] incurs significantly higher computational cost as it aligns gradients across all parameters. Notably, for the largest dataset, ShapeNet [4], DC requires over 7 hours of training, whereas DM completes training in just 9 minutes (0.15 hours), making DC more than 50 times slower. This latency is even more pronounced in point cloud datasets, as networks designed for point cloud processing, such as PointNet [14], have higher computational complexity than those used in image dataset distillation. Since our algorithm is based on distribution matching approach, it requires only a marginal increase in training time compared to the DM algorithm, yet achieves a significant improvement in performance.
Layer 1 Layer 2 Layer 3 PointNet
_ _ 61.47
_ _ 62.22
_ _ 62.72
Table 7: Ablation study of the impact of different PointNet layers. The performance was averaged across the four models of ModelNet10 [24], ModelNet40 [24], ShapeNet [4], and ScanObjectNN [19] datasets. Experiments were performed at PPC 3.
4.3 Visualization
Figure 3 presents a visualization of the distilled synthetic dataset generated by previous methods and the proposed method. DC [29], which relies on gradient matching, struggles to effectively match the gradients of the large PointNet [14] network. As a result, it fails to deviate significantly from the initialized point cloud model, leading to minimal movement in certain points while introducing noise. DM [28], on the other hand, performs feature distribution matching after the global max pooling operation in PointNet, which loses critical information. As a result, similar to DC, DM preserves the original shape while only shifting specific points, failing to capture meaningful structural changes. In contrast, the proposed method successfully preserves the semantic structure of each class while selectively learning essential features. Notably, in the airplane class, the wing structure appears distinct compared to other methods. Additionally, in certain categories, such as guitar, both the shape and orientation are changed, demonstrating that the proposed method functions as intended. These results highlight the effectiveness of our approach in producing a cleaner, noise-free synthetic dataset with superior visual quality compared to previous methods.
4.4 Ablation Study
4.4.1 Effect of Permutation Invariant Distribution Matching Loss
Table 5 presents an ablation study on the impact of the proposed sorting mechanism on classification performance using PointNet [14]. Without the sorting mechanism (w/o $\mathcal{F}$), the unordered nature of point clouds restricts effective feature matching, resulting in performance that is nearly identical to that of random selection [16], indicating that the model has not been properly optimized. Otherwise, when the sorting mechanism is applied (w/ $\mathcal{F}$), accuracy improves across all datasets, as the sorted feature provide a consistent and structured representation that facilitates more effective feature matching during optimization. This result confirms that addressing the inherent permutation-invariance of point clouds through sorting enables the network to align corresponding features more effectively, resulting in a substantial improvement in classification performance.
4.4.2 Effect of Orientation Optimization
Table 6 presents an ablation study to evaluate the impact of optimizing the synthetic dataset and rotation angles independently and jointly, using the proposed PIDM loss. We evaluated on dataset groups categorized into aligned, mixed, and rotated to analyze the effectiveness of the proposed orientation optimization. The aligned group consists of datasets where objects maintain a consistent orientation, including ModelNet10 and ShapeNet [4]. The mixed group includes datasets where only certain classes exhibit rotation variations, represented by ModelNet40 [24]. The rotated group comprises datasets in which objects exhibit arbitrary rotations across all classes, such as ScanObjectNN [19].
In the second row, we present the results where only the rotation angle is optimized without optimizing the synthetic dataset. Introducing the proposed orientation optimization method, optimized using the PIDM loss, improves performance across all groups compared to the non-optimized setting. The improvement becomes more noticeable as the class with rotations becomes more dominant in the dataset, indicating the importance of orientation optimization in determining the optimal direction for data with arbitrary rotations. The addition of synthetic dataset optimization further boosts performance, with the best results achieved when both optimizations are applied jointly. Specifically, the accuracy on the rotated dataset increases significantly, showing that jointly optimizing the synthetic dataset and rotation angles with the PIDM loss helps the model better adapt to rotation variations.
4.4.3 Effect of Feature Extraction Layers in PointNet
Table 7 presents an ablation study examining how classification performance on PointNet [14] varies when using feature maps from different layers. PointNet consists of three layers before the pooling operation, which produce feature maps with 64, 128, and 1024 channels, respectively. The results show that deeper feature extraction layers generally lead to better performance. This performance improvement can be attributed to the increased number of channels, which allows the feature map to capture more detailed and diverse information. A higher-dimensional representation provides greater capacity to maintain essential features from the original dataset, enabling the network to better approximate the underlying distribution.
5 Conclusion
We proposed a permutation-invariant and orientation-aware dataset distillation framework for 3D point clouds. To alleviate the inconsistency of point ordering between compared 3D models, we devise a Permutation Invariant Distribution Matching (PIDM) loss where the sorted features are compared. We also employed learnable parameters of rotation angles to estimate optimal orientation of synthetic models. The geometric structures and orientation of the synthetic models are jointly optimized during the dataset distillation. Experimental results evaluated on four widely used benchmark datasets of ModelNet10 [24], ModelNet40 [24], ShapeNet [4], and ScanObjectNN [19] demonstrate that the proposed method outperforms existing distillation methods while maintaining strong cross-architecture generalization. Furthermore, we also validated the effectiveness of each component of the proposed method through extensive ablation studies.